dont know if im in the right place direct me if im not but.. wut is
At a pet store, Rhonda paid $11.50 for 3 dog bowls and 4 bones. Kelly paid $13 for 2 dog bowls and 8 bones. How much will Andrew pay for 4 dog bowls and 2 dog bowls
this is wut i got so far
$11.50 = 3 dog bowls + 4 bones
$13 = 2 dog bowls + 8 bones
? = 4 dog bones + 2 bones
walk thrue if you can plz

thank you ahead of time

Answer by checkley77(12844)

$11.50 = 3 dog bowls + 4 bones
$13 = 2 dog bowls + 8 bones
11.50=3d+4b multiply this equation by -2 & add.
13=2d+8b
-23=-6d-8b
--------------
-10=-4d
-4d=-10
d=-10/-4
d=$2.50 the price of one dog bowl.
11.50=3*2.50+4b
11.50=7.50+4b
4b=11.50-7.50
4b=4
b=$1 the price of one bone.
4d+2b=4*2.50+2*1=10+2=$12 for 4 dog bowls plus 2 bones.